THE b ^{3}Π_{u0} STATE OF ^{39}K_{2}

\underline{A. J. ROSS}, F. MARTIN, P. CROZET, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France; A. M. LYYRA, Department of Physics, Temple University, Philadelphia, PA 19122-6082, U.S.A; LI LI, Department of Modern Applied Physics, Tsinghua University, Beijing 10084 China; T. BERGEMAN, Department of Physics and Astronomy, State University of New York, Stony Brook, NY 11794-3800 USA; C. AMIOT, Laboratoire Aimé Cotton (CNRS UPR3321), Campus d’Orsay, Bâtiment 505, 91405 Orsay, France; M. RIAD MANAA, Energetic Materials Center, Lawrence Livermore National Laboratory, University of California, Livermore, CA 94551 USA.

New observations of the b ^{3}Π_{u0} state of K_{2} have been made by laser induced fluorescence spectroscopy, following single colour OODR excitation with a cw Ti:sapphire laser. Fourier transform spectra revealed two pieces of a progression of the 2 ^{3}Π_{g} → b ^{3}Π_{u0} system, beginning around 14250 cm^{−1}, at v = 0, and terminating at 7200 cm^{−1}. The two pieces were unfortunately separated by very strong A → X emission, which dominated the region 8000 - 12000 cm^{−1}, so that no measurements could be made for many intermediate vibrational levels. The triplet emission in the visible region was severely affected by spin-orbit interactions with A ^{1}Σ_{u}^{+}, whilst the infrared part of the spectrum was much less perturbed. The intermediate part of the potential is known only through the interactions of the A ^{1}Σ_{u}^{+} and b ^{3}Π_{u} states. The spin-orbit interaction between the A and b states up to v_{b} = 90 has been modelled using an ab initio calculation of the spin-orbit function, which shows a notable dip as a function of R. This model allowed us to assign the vibrational quantum numbers in the infrared part of the fluorescence spectrum : the observations cover 89 ≤ v ≤ 118. We are fitting the available A and b state data to Y_{ij} parameters. At present, we have achieved fits (with a standard deviation of 0.025 cm^{−1}) of about 1000 term values or combination differences, corresponding to levels up to 5500 cm^{−1} above the minimum of the b state. We intend to continue the fits over a further 1500 cm^{−1} to incorporate all the data, reaching term values 500 cm^{−1} below the 4S + 4P threshold. An asymptotic calculation of the Hund’s case c) 0_{u}^{+} states should be able to extend a deperturbed potential curve for the b state sufficiently accurately to allow the assignment of existing photoassociation data.